OpenStudy (anonymous):

Brad invested $3,200 in a five year CD that pays three percent compounded annually. What is the compound interest and the amount that will be in the bank after three years?

OpenStudy (wolf1728):

Total = Principal * (1 + rate) ^ years

OpenStudy (wolf1728):

Total = 3,200 * (1.03)^3 (we want the 3 year total) Total = 3,200 * 1.092727 Total = 3,496.7264 Total = 3,496.73

OpenStudy (wolf1728):

Everyone agree?

OpenStudy (anonymous):

3;200(1.033) 218.5454 . 3,200 * 218.5454 = 3;(200)(218.5454) 43709.08

OpenStudy (anonymous):

am i wrong or right

OpenStudy (wolf1728):

3,200 + 218.5454 = 3,418.5454

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that correct or wrong

OpenStudy (anonymous):

do you have option on this question

OpenStudy (wolf1728):

Wow I got thrown off Open Study No, your answer is wrong

OpenStudy (wolf1728):

Basically, you have to calculate 3,200 * (1.03)^3 Do you know how to do that?

OpenStudy (anonymous):

yes